# Data-Driven Prediction and Uncertainty Quantification of PWR Crud-Induced Power Shift Using Convolutional Neural Networks

### v1

@@ 1 Introduction
The organization of the subsequent sections in this paper is as follows: Section 2 introduces the problem and details the data employed. Section 3 outlines the general approach, architecture of the neural network used, and UQ methodology. In Section 4, we present the findings of our investigation, encompassing the CIPS predictions and the associated UQ. Finally, our concluding remarks on this study can be found in Section 5.
@@ 2.2 ML Model for CIPS Prediction
CNNs have gained widespread adoption in computer vision approaches, with applications such as handwriting detection and object recognition. Originally based on the feline visual cortex, convolutional layers allow for pattern recognition using the geometric similarity between relative entry locations within data rather than absolute [11][12]. CNNs have also gained attention in the field of nuclear engineering, with applications such as fault diagnosis [13][14]. This technique offers significant benefits such as the preservation of spatial information, hierarchical feature learning, and translational invariance [15]. In addition, these networks are more efficient (requiring fewer parameters) compared to a similarly-tasked fully-connected network. These characteristics make CNNs superior when dealing with data exhibiting spatial correlations, making them appropriate to use with the 3D pin powers.
Acknowledging the existence of inherent uncertainties, we intend to provide insight into the confidence levels of the ML model’s predictions through an UQ method. Specifically, we will employ the MCD-based UQ technique, known for its effectiveness in quantifying uncertainties of the ML models.
@@ 3.1 General Approach
To illustrate the general process of this study, a visualization of the workflow is provided in Figure 2. The core model data, generated using MPACT, are the inputs to the network and consist of the 3D pin powers, core exposures, and boron concentrations. For each of these entries, the corresponding CIPS Index calculated based on plant measurement data is then added. These arrays for Cycles $7$ and $8$ are then combined, including relevant information to identify the entries later on, such as assembly ID and cycle number. During pre-processing, the CIPS classes are first computed by setting a hard threshold of $<10$ in Index for the positive class and $\leq 10$ for the negative class. The resulting vector of positive and negatives (binary encoded) was then added to the data array. This is then scaled using min-max scaling, shuffled, and partitioned using an $80$%/$10$%/$10$% split for training, validation, and testing of the ML model. This data is then fed into the hyperparameter optimization algorithm, which trains a set number of models with unique hyperparameter configurations to minimize the validation loss. Once this process was complete, the best-performing model’s hyperparameters were set for use in $k$-fold cross validation (CV).
@@ 3.3.1 Overview of MCD for UQ of DNN
Typical DNNs are known as “black box” models due to their inherent complexity, making clear interpretation difficult. The technique of MCD has emerged as an approach to help mitigate these concerns in the field of UQ as a Bayesian approximation [29]. The basis of this approach is in leveraging the random nature of dropout layers in selectively disabling neurons for a given pass through the network, as shown in Figure 4. Typically, these layers are only active during the training step, but in MCD they are allowed to be active in the prediction step as well after the training is completed. By introducing active dropout during both training and prediction, MCD allows for an arbitrary number of predictions to be generated for a single input. The collection of predictions for a given input can be interpreted as random samples of the output distribution, permitting statistical analysis beyond what conventional DNN outputs are capable of. This technique also has the potential to reduce the risk of overfitting, particularly in cases of noisy or uncertain data [30]. With the use of MCD, it is necessary to determine a set of effective uncertainty metrics to assess the performance of the model within this lens [31].
@@ 4.1 Results for Base Model
Each of the five folds was trained over a period of $100$ epochs using their unique training datasets. On average, each fold took $43.74$ minutes to train using four Apple M1 Pro cores in resources. The five folds, which are uniquely-weighted networks, were then frozen and used to predict against their test sets. The combined inference time of these folds was $0.0172$ seconds to essentially predict two cycles worth of data. The breakdown of these time expenses is located in Table 3. The lightweight nature of the trained model is one of the primary benefits of using neural networks. In the case of this study, the total training time of $215.9$ minutes while using limited resources would likely not be prohibitive to its deployment, as the network could be re-trained in relatively short time in situations where more data becomes available to update the model.
@@ 4.2 Results for Uncertainty Quantification
To implement MCD, the same methodology utilizing $k$-fold CV was maintained from Section 3.1. The training process was identical to that of the base model, but once the network was trained, the predictions made on the test cases were subject to MCD. This was performed by setting the model to prediction-mode while also allowing it to be trainable. The weights of this model are still frozen, but it allows for the dropout layer parameters to vary stochastically. The impact of this is that when the same input enters the neural network model multiple times, different outputs will be obtained since different neurons are dropped each time, the ensemble of which can represent the uncertainty in the output.
@@ 4.2.1 MCD Prediction Performance
Over the course of the $5$ folds, each of the test inputs was predicted for a total of $T=50$ times to produce $50$ distinct output values while using a dropout rate of $0.1$. For each of these sets of $50$, the means and variances were then taken for an assessment of their uncertainties. The means were then treated in the same manner as the base model’s output probabilities, and were split at a threshold of $0.5$ to determine the predicted CIPS class. With the vector of class predictions and the reference values, the conventional and class-balanced metrics were then applied and reported in Table 6. When compared to the performance results of Section 4.1.1, the MCD results achieve the same level of accuracy while exhibiting slightly reduced performance in other metrics. The differences in these metrics with those of the base model are less than $0.01$ in all cases. Although the technique of MCD does not increase the performance of this classifier within these metrics, the close agreement of these values and those of the base model reflects the classifier’s ability to consistently and repeatably make predictions [37].
@@ 5 Conclusions
With the results enclosed in this study, this approach has the potential to be useful in a core design setting for the rapid prototyping of loading patterns by providing information about potential CIPS effects. A mature crud model could be feasibly included in a full-core ML model [43]. Future work would be focused on improving the performance of the method, investigating the generalization to other reactors, and a more extensive validation process to ensure the trustworthiness and reliability of such a model. To accomplish these, a larger set of data would be required, ideally from multiple reactors. Transitioning analysis from MPACT-based data to a core simulator commonly used in industry calculations would also be necessary to further develop this approach into a potential core design tool. To explore the applicability of a network trained on one reactor’s data being applied to another reactor, the use of techniques such as transfer learning may be helpful in a fine-tuning regime. Overall, there are several aspects of this method that could be further improved and investigated to create a robust and validated framework.
Acknowledgement

### v2

@@ 1 Introduction
The organization of the subsequent sections in this paper is as follows: Section 2 introduces the problem and details the data employed. Section 3 outlines the general approach, architecture of the neural network used, and UQ methodology. In Section 4, we present the findings of our investigation, encompassing the CIPS predictions and the associated UQ. Discussion about the practical implementation and future work is then provided in Section 5. Finally, our concluding remarks on this study can be found in Section 6.
@@ 2.2 ML Model for CIPS Prediction
CNNs have gained widespread adoption in computer vision approaches, with applications such as handwriting detection and object recognition. Originally based on the feline visual cortex, convolutional layers allow for pattern recognition using the geometric similarity between relative entry locations within data rather than absolute [11] [12]. CNNs have also gained wide attention in the energy domain. For example, in nuclear energy, CNN-based fault analysis and diagnosis models have been developed for large commercial nuclear power plants [13] [14], as well as small modular reactors [15]. Stefenon et al. [16] developed a hyper-tuned wavelet CNN with long short-term memory for time series forecasting in hydroelectric power plants. In the field of load prediction or forecasting, a multi-scale fusion CNN was presented in [17] for load prediction of integrated energy systems. Further, Xu et al. [18] added attention mechanism in a time series depthwise separable CNN in a framework for electricity load forecasting. The CNN model offers significant benefits such as the preservation of spatial information, hierarchical feature learning, and translational invariance [19]. In addition, these networks are more efficient (requiring fewer parameters) compared to a similarly-tasked fully-connected network. These characteristics make CNNs superior when dealing with data exhibiting spatial correlations, making them appropriate to use with the 3D pin powers.
Acknowledging the existence of inherent uncertainties, the novelty of this work is to provide insight into the confidence levels of the CNN ML model’s predictions through an UQ method, when the training dataset is relatively limited in size. Specifically, we will employ the MCD-based UQ technique, known for its effectiveness in quantifying uncertainties of the ML models.
@@ 3.1 General Approach
To illustrate the general process of this study, a visualization of the workflow is provided in Figure 2. The core model data, generated using MPACT, are the inputs to the network and consist of the 3D pin powers, core exposures, and boron concentrations. For each of these entries, the corresponding CIPS Index calculated based on plant measurement data is then added. These arrays for Cycles $7$ and $8$ are then combined, including relevant information to identify the entries later on, such as assembly ID and cycle number. During pre-processing, the CIPS classes were computed by setting a hard threshold: an index value above 10 was classified as positive, and 10 or below as negative. This resulted in a binary-encoded vector of positive and negative labels. Each input channel (pin powers, core exposures, and boron concentrations) was independently scaled using min-max scaling (range: [0, 1]), shuffled, and partitioned using an $80$%/$10$%/$10$% split for training, validation, and testing of the ML model. Scaling ensures that the model does not develop biases due to differences in magnitude, and shuffling prevents bias due to the ordering of instances. This data is then fed into the hyperparameter optimization algorithm, which trains a set number of models with unique hyperparameter configurations to minimize the validation loss. Once this process was complete, the best-performing model’s hyperparameters were set for use in $k$-fold cross validation (CV).
@@
The prevention and detection of overfitting, defined as a decline in a model’s ability to generalize to new data despite high performance on training data, is an important concern in this study. The use of two cycles, totaling 2352 input sets with a class imbalance (15.2% of input pairs qualify as positive CIPS), poses a potential risk of overfitting. To mitigate this, careful architecture design and hyperparameter tuning, as described above, balance model complexity to ensure the necessary features are learned without memorizing the input data. Exponential learning rate decay is employed to gradually reduce weight update rates, ensuring smooth convergence to the optimal solution. Training and validation losses are monitored to detect discrepancies that may indicate overfitting. Finally, the $k$-fold CV methodology ensures that the model generalizes well across all sections of the data, a strong indicator of strong performance on unseen datasets.
@@ 3.3.1 Overview of MCD for UQ of DNN
Typical DNNs are known as “black box” models due to their inherent complexity, making clear interpretation difficult. The technique of MCD has emerged as an approach to help mitigate these concerns in the field of UQ as a Bayesian approximation [33]. The basis of this approach is in leveraging the random nature of dropout layers in selectively disabling neurons for a given pass through the network, as shown in Figure 4. Typically, these layers are only active during the training step, but in MCD they are allowed to be active in the prediction step as well after the training is completed. By introducing active dropout during both training and prediction, MCD allows for an arbitrary number of predictions to be generated for a single input. The collection of predictions for a given input can be interpreted as random samples of the output distribution, permitting statistical analysis beyond what conventional DNN outputs are capable of. This technique also has the potential to reduce the risk of overfitting, particularly in cases of noisy or uncertain data [34]. MCD, like all methods, is not immune from drawbacks. Its sensitivity to the dropout rate, which determines how many neurons are inactive, can greatly influence both a model’s performance and uncertainty estimates [35]. Higher rates may increase the regularization too much, leading to underfitting via excessive information loss, while rates too small can lead to overly confident and poorly calibrated uncertainties. Another drawback is the associated computational overhead, which can be considerable in resource-intensive settings, since inference requires multiple forward passes. MCD can also be less effective for tasks using recurrent neural networks, long short-term memory architectures, or structured outputs like image segmentation. However, for two-class problems, such as in this study, MCD is well-suited since the output probability can be modeled as a distribution with a mean and variance. Finally, to effectively use MCD in this context, defining appropriate uncertainty metrics is crucial for evaluating model performance [36].
@@ 4.1 Results for Base Model
Each of the five folds was trained over a period of $100$ epochs using their unique training datasets. On average, each fold took $43.74$ minutes to train using four Apple M1 Pro cores in resources. The five folds, which are uniquely-weighted networks, were then frozen and used to predict against their test sets. The combined inference time of these folds was $0.0172$ seconds to essentially predict two cycles worth of data. The breakdown of these time expenses is located in Table 3.
The lightweight nature of the trained model is a significant advantage. By comparison, the current coupled CTF/MAMBA methodology, integrated with the SIMULATE core simulator, typically requires one to two hours of runtime using 193 cores [41]. Although the MAMBA approach offers higher-fidelity outputs, such as pinwise crud mass projections, the neural network in this work is designed for the rapid prototyping of core loading patterns. Finalized designs would then be evaluated using the traditional methods that are currently in use. With entire-core, full-cycle prediction times under 20 milliseconds, the model could be readily incorporated into existing drag-and-place graphical user interface tools for first-order LP prototyping.
The total training time of $215.9$ minutes using limited resources is unlikely to hinder deployment. This method’s small computational needs would allow the model to be re-trained quickly when new data becomes available, building a more effective model over time. Overall, these time savings when compared to traditional physics-based models make it a viable tool to improve workflow efficiency in the core design process.
@@ 4.2 Results for Uncertainty Quantification
To implement MCD, the same methodology utilizing $k$-fold CV was maintained from Section 3.1. The architecture was modified to apply dropout on all hidden layers, to ensure that uncertainty is propagated throughout the network. It is of note that three-dimensional spatial dropout was applied to the convolutional elements of the model, as traditional dropout can remove features that may be spatially important [43]. The training process was identical to that of the base model, but once the network was trained, the predictions made on the test cases were subject to MCD. This was performed by setting the model to prediction-mode while also allowing it to be trainable. The weights of this model are still frozen, but it allows for the dropout layer parameters to vary stochastically. The impact of this is that when the same input enters the neural network model multiple times, different outputs will be obtained since different neurons are dropped each time, the ensemble of which can represent the uncertainty in the output.
@@ 4.2.1 MCD Prediction Performance
Over the course of the $5$ folds, each of the test inputs was predicted for a total of $T=50$ times to produce $50$ distinct output values while using a dropout rate of $0.1$. This value was empirically optimized using a grid search with the validation partition, with the objective to maximize prediction performance while maintaining reasonable uncertainty estimates (evaluated via their calibration). For each of these sets of $50$, the means and variances were then taken for an assessment of their uncertainties. The means were then treated in the same manner as the base model’s output probabilities, and were split at a threshold of $0.5$ to determine the predicted CIPS class. With the vector of class predictions and the reference values, the conventional and class-balanced metrics were then applied and reported in Table 6. When compared to the performance results of Section 4.1.1, the MCD results achieve the same level of accuracy while exhibiting slightly reduced performance in other metrics. The differences in these metrics with those of the base model are less than $0.01$ in all cases. Although the technique of MCD does not increase the performance of this classifier within these metrics, the close agreement of these values and those of the base model reflects the classifier’s ability to consistently and repeatably make predictions [44].
@@ 4.2.2 Model Calibration
As a quantitative assessment of calibration, the non-MCD model achieved a Brier score of 0.054, while the MCD model reported a slightly improved score of 0.052. The Brier score, which measures the mean squared difference between predicted probabilities and actual outcomes, ranges from 0 (perfect calibration and accuracy) to 1 (completely miscalibrated or inaccurate predictions). Although the improvement in the Brier score appears modest, it reflects a measurable improvement in the MCD model’s ability to provide more accurate and calibrated probability estimates. This reduction indicates that the MCD model better aligns its predicted probabilities with the true likelihood of outcomes, particularly in regions where the non-MCD model exhibited overconfidence or underconfidence in its predictions. When combined with visual evidence from the reliability diagrams, these results suggest that MCD not only improves calibration but also improves the overall reliability of the model’s probabilistic outputs.
@@ 4.2.4 Bivariate Analysis
5 Discussion
The above results demonstrate that a surrogate model can effectively predict instances of CIPS in CNS-1 by using a combination of core model and measurement data from previous cycles. When considering the deployment of this approach to an operational capacity in the frame of industry, practical concerns must be addressed. The most important is defining the scope of its use case, which also is relevant from a regulatory standpoint.
It is unlikely that such a framework would be of use as a means of evaluating a finalized core design, since utilities and vendors have established methods that are well-accepted in the regulatory environment. This method’s use, however, is to instead be a first-order estimation tool to better-inform core designers while they are in the phase of prototyping loading patterns. This process often requires several potential designs to be run, and with the current physics-based crud tools, can take a significant amount of time per design as noted in Section 4.1. Rather than each prospective design being assessed with these conventional tools, the approach described in this study would effectively eliminate that time cost by providing accurate real-time estimates even if they are lower-fidelity. By being involved in optimization step, with a finalized design still being evaluated by conventional means, it would not be involved in any final safety-related workflow processes or formal analyses, a benefit from a regulatory standpoint.
From a practical standpoint, a well-trained model could be implemented into a nodal core simulator [50], many of which have a graphical user interface to place assemblies around a quarter or eighth-core to obtain a heat map of information like power peaking factors. With the CIPS model’s inferences made on the scale of milliseconds, it would be fast enough to provide CIPS potential on an assembly-wise basis as the assemblies are being configured. A potential visualization of this would be a heat map overlay such as the visualization in Figure 9 of Section 4.2.3. Not only would this provide the predicted probability of each assembly’s CIPS state, but it would also provide the predictions’ confidences to help better-inform the core designer.
In terms of progressing a model built on this methodology to a mature state, substantive validation is required not only at CNS-1 but to also determine if it generalizes to other reactors. This specific model’s ability to do this is unknown, as each reactor’s CIPS behavior is typically unique. The model in this study is specifically calibrated to CNS-1 to support the objective to construct and validate the methodology itself. Improving performance and determining its generalization would require more data from CIPS-affected cycles in different reactors. The data problem is most challenging aspect of implementing this approach, as both operating data and core design data are difficult to obtain. Accomplishing this would likely require partnership with a utility and/or reactor vendor with access to large amounts of this data. Once data is obtained, however, the model can be rapidly updated to improve its performance and reliability on specific reactor units, as well as in exploring the transferability of shared features between different units with techniques such as transfer learning.
@@ 6 Conclusions
Future work was also discussed, with particular emphasis on the steps necessary for such a method to be deployed to assist in core design in both CNS-1 and new reactors. With the results enclosed in this study, this approach has the potential to be useful in a core design setting for the rapid prototyping of loading patterns by providing information about potential CIPS effects.
Acknowledgment